# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v5

@@ 5.5 Anchoring to Teacher Policies: Continuous Control Distillation
We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence. For each environment, we collect a distillation dataset of 10K states, each associated with $K=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%). Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts. Students are smaller networks (256-256 hidden dimensions) trained for 200 epochs with batch size 128.
@@
  • ADPO-Teacher-Anchor: Student anchored to frozen teacher policy $\pi_{\mathrm{teacher}}$, enabling both preference learning and geometric regularization.
Table 6: Policy distillation on continuous control (MuJoCo). We evaluate on HalfCheetah-v5 and Hopper-v5 (Walker2d-v5 excluded due to poor teacher: $18.86\pm 5.75$). Results averaged over 5 seeds per environment. ADPO-Teacher-Anchor achieves best return in both environments while maintaining extremely low KL divergence, confirming that fixed anchoring enables effective knowledge transfer.
@@
HalfCheetah-v5 (Teacher: 53.5) KD $311.2\pm 10.4$ 0.984 23.73
ADPO-Self-Anchor $131.7\pm 23.3$ 0.626 19.27
ADPO-Teacher-Anchor $\mathbf{206.7\pm 24.7}$ 0.667 0.08
Hopper-v5 (Teacher: 107.5) KD $25.1\pm 5.9$ 0.990 52.78
ADPO-Self-Anchor $34.1\pm 4.5$ 0.587 81.30
ADPO-Teacher-Anchor $\mathbf{65.4\pm 1.7}$ 0.769 0.01
@@
Table 6 validates our hypothesis that fixed anchors excel at offline distillation:
@@
    ADPO-Teacher-Anchor dominates on both environments. On HalfCheetah-v5, it achieves 206.7 (387% of teacher’s 53.5), significantly outperforming KD’s 311.2-311.2 despite KD’s near-perfect ranking quality (NDCG=0.984). On Hopper-v5, ADPO-Teacher-Anchor reaches 65.4 (61% of teacher’s 107.5), again surpassing KD’s 25.1 (23%). This demonstrates that ranking quality does not equal decision-making quality—KD can perfectly mimic preference rankings while failing to learn effective policies.
@@
    ADPO-Self-Anchor underperforms (131.7-131.7 and $34.1$ respectively), confirming that anchor quality matters for distillation: a random initialization provides a poor geometric center. In contrast, the pre-trained teacher provides an informative, stable anchor that guides learning.
@@
    Task-dependent anchor selection validated. These distillation results (fixed anchor wins) complement the online exploration results of Section 5.4 (dynamic anchor wins), confirming our hypothesis:
@@
    • Offline distillation $\to$ fixed anchor provides stable teacher targets
@@
These results directly validate our geometric framework (Section 3, Lemma 3.5):
  • KD vs. ADPO: KD operates in “absolute probability space” without a fixed geometric center (Table 8). ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@
Practical implications.
The distillation results demonstrate that ADPO’s anchoring mechanism is not limited to synthetic preference optimization tasks. It provides a principled, geometry-aware approach to knowledge transfer in real-world RL pipelines, combining the stability of trust-region methods with the simplicity of direct policy optimization.
(a) Return comparison across environments
(b) KL divergence vs Return trade-off
Figure 4: Policy distillation results on MuJoCo environments. Teacher performance: HalfCheetah-v5 (53.5), Hopper-v5 (107.5). Left: Student return comparison across both environments. Right: Trade-off between return and KL divergence to teacher. ADPO-Teacher-Anchor achieves best return in both environments (206.7 on HalfCheetah = 387% of teacher; 65.4 on Hopper = 61% of teacher) while maintaining extremely low KL (0.08 and 0.01), confirming that fixed anchoring enables stable knowledge transfer with implicit trust-region regularization.
5.6 Hyperparameter Sensitivity
To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$. Figure˜5 shows the results. Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$. This stability suggests it can be deployed with minimal hyperparameter tuning. In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
(a) Pairwise anchored
(b) Listwise raw
(c) Listwise KDE
Figure 5: Temperature sensitivity ablation. Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment. KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
6 Discussion
Why does anchoring work?
Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$. The crucial difference: Standard DPO anchors the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$, while ADPO anchors each score individually before comparing. This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.5)—anchoring is trust region by design. Our experiments demonstrate relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor.
Hard vs. soft labels.
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that "average out" signal. However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence. Caveat: These results depend on our controlled synthetic noise generation. In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
Listwise dominance.
Listwise methods win in 9/12 scenarios (Table 3) by leveraging full group information and anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases—especially valuable when rewards have group-level shifts (different annotators with different baselines).
@@ 6 Discussion
Knowledge Distillation ADPO-Teacher-Anchor Superior returns & stability (e.g., +206 vs -311)
@@ 8 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms. By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(s-s^{\mathrm{ref}})/\tau$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature. This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
@@ D.1 Noise Generation Procedures
  • Light: SNR = $\frac{\operatorname{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
  • Medium: SNR = 0.5, $p_{\text{out}}=0.10$
  • Heavy: SNR = 0.2, $p_{\text{out}}=0.20$
@@ E.3 Temperature Sensitivity (Full Grid)
Figure 5 in the main text shows three representative methods. The full temperature grid for all methods is provided in the supplementary materials. Key findings:

### v6

@@ 5.5 Anchoring to Teacher Policies: Continuous Control Distillation
We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence. For each environment, we collect a distillation dataset of 10K states, each associated with $P=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%). Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts. Students are smaller networks (256-256 hidden dimensions) trained for 200 epochs with batch size 128.
@@
  • ADPO-Self-Anchor-EMA: Student anchored to an exponential moving average (EMA) of its own policy, providing a slowly moving reference that tightens coupling while remaining stable.
Table 5: Policy distillation on continuous control (MuJoCo). Results averaged over 5 seeds per environment. ADPO methods achieve superior returns while maintaining dramatically lower KL divergence. ADPO-self-anchor uses fixed initial policy as anchor; ADPO-self-anchor-EMA uses exponential moving average. Note: Anchor strategy shows task-dependent effectiveness—fixed anchor excels on HalfCheetah, moving anchor on Hopper.
@@
HalfCheetah-v5 (Teacher: 8476.5) KD $309.0\pm 98.2$ 0.857 30.50
ADPO-self-anchor $\mathbf{279.3\pm 53.2}$ 0.765 10.45
ADPO-self-anchor-EMA $166.8\pm 227.0$ 0.858 0.62
Hopper-v5 (Teacher: 1169.8) KD $36.9\pm 20.2$ 0.799 16.98
ADPO-self-anchor $177.3\pm 158.4$ 0.684 7.61
ADPO-self-anchor-EMA $\mathbf{855.6\pm 269.1}$ 0.755 3.97
@@
Table 5 validates our hypothesis that anchoring excels at offline distillation, with task-dependent anchor strategies:
@@
    ADPO dramatically outperforms KD on both environments. On HalfCheetah-v5, ADPO-self-anchor achieves 279.3 vs KD’s 309.0-309.0 (588-point improvement), while on Hopper-v5, ADPO-self-anchor-EMA reaches 855.6 (73% of teacher’s 1169.8) vs KD’s 36.9 (3%). This demonstrates that ranking quality does not equal decision-making quality—KD achieves high NDCG (0.857, 0.799) but fails to learn effective policies.
@@
    Task-dependent anchor strategy confirmed. Fixed anchor (ADPO-self-anchor) achieves best return on HalfCheetah (279.3), while moving anchor (ADPO-self-anchor-EMA) excels on Hopper (855.6). Both achieve lower KL than KD, with moving anchor providing the tightest coupling (0.62 and 3.97). This validates the theoretical prediction from Section 3.5.
@@
    Anchor strategy trade-off validated. These distillation results complement the online exploration results of Section 5.4, confirming our hypothesis:
@@
    • Offline distillation $\to$ both fixed and moving anchors excel, with task-dependent preferences
@@
These results directly validate our geometric framework (Section 3.2, Lemma 3.6):
  • KD vs. ADPO: KD operates in absolute probability space without a fixed geometric center. ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
  • Anchor flexibility: Whether anchoring to an old policy (online RL) or to a fixed reference (e.g., the initial policy or an EMA of the student in offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@
Knowledge Distillation ADPO-Self-Anchor Superior returns & stability (e.g., 855.6 vs 36.9 on Hopper; 279.3 vs -309.0 on HalfCheetah)
@@ 7 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms. By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature. This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
@@ D.1 Noise Generation Procedures
  • Light: noise scale $\sigma=1.0$, $p_{\text{out}}=0.05$
  • Medium: noise scale $\sigma=2.0$, $p_{\text{out}}=0.10$
  • Heavy: noise scale $\sigma=2.5$, $p_{\text{out}}=0.20$
@@ E.3 Temperature Sensitivity (Full Grid)
Temperature sensitivity analysis shows that ADPO maintains robust performance across a wide range of temperature values. Key findings: